OpenStudy (anonymous):

Solve the system: 22x-19y=28. 55x-29y=107.

OpenStudy (amistre64):

cramer!!

OpenStudy (amistre64):

22x-19y=28 55x-29y=107 y = \(\cfrac{107(22)-55(28)}{-29(22) - 55(-19)}\) x = \(\cfrac{107(-19)-28(-29)}{55(-19)- 22(-29) }\)